an element crystalises in fcc and has edge length 0.56 nm. calculate the density of the element molar mass of the element is 40g mol inverse

Grade:12

1 Answers

Sanskar
one year ago

The density of the crystal is given by the formula:

d = ZM/a^3NA

Given data Z for FCC = 4; M (atomic mass) = 40 g/mol; a (inter-atomic distance) = 0.556 nm or 0.556x10^-9 m, Substituting the values in the formula the density would be

 d = (4*40 g)/(0.556x10^-9m)^3*6.023x10^23

  = 160/(0.1719x10^-27m3) *6.023x10^23

  = 1.546x10^6 g/m3

  d = 1.546 g/cm3